OpenStudy (anonymous):

19. If the area of a circle is 8 pi , what is the length of its diameter?

OpenStudy (king):

\[2\sqrt{8}\]

OpenStudy (king):

or \[\sqrt{32}\]

OpenStudy (anonymous):

it's either 4 8 2√2 4√2

OpenStudy (anonymous):

no, its 4

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

it wasn't 4

OpenStudy (anonymous):

sorry im a dumbass haha its 4 root 2

OpenStudy (anonymous):

last one, thats my bad

OpenStudy (katrinakaif):

a = pie r ^2 8 pi = pi r^2 8 = r^2 r = 2 square root 4 Now double it.

OpenStudy (king):

Thats what i said 4 root 2 is the same as 2root 8 or root 32

OpenStudy (katrinakaif):

yes its what you said. Never said you where incorrect.